A racing car moving along a circular track of radius 'r' at constant speed 'v' has described the angle 'a' at the centre of track. What is the average velocity?
1
Expert's answer
2020-12-21T11:01:16-0500

Answer

Total displacement =a

Total time taken to travell this displacement

"T=\frac {\pi a }{3v}"


So average velocity

"v_a=\frac{displacement}{time}"


average velocity

"=\frac{3v}{\pi}"